Use the pythagorean theorem then find the exact length of the hypotenuse in the following right triangle
5 and 7

Answer by stanbon(75887)

Use the pythagorean theorem then find the exact length of the hypotenuse in the following right triangle
5 and 7
-----------------------
hypoteuse^2 = 5^2 + 7^2
h^2 = 25+49
h^2 = 74
hypotenuse = sqrt(74)
===========================
Cheers,
Stan H.